OpenStudy (anonymous):

In the equation f(x)=7, using the formula f(x+h)-f(x)/h, how do they get 0. How does f(x+h)=0 also?

OpenStudy (anonymous):

I mean =7

OpenStudy (anonymous):

i would think, by h =0 is the only way to get f(x+h) = 7

OpenStudy (anonymous):

how do we know h=0?

OpenStudy (anonymous):

Okay, I understand now. How do we get rid of the bottom h then?

OpenStudy (anonymous):

h is still zero

OpenStudy (anonymous):

but if that would leave the equation 0/0 which is undefine? How can i get rid of the h?

OpenStudy (anonymous):

your function is a constant, so \(f(x)=7\) and also \(f(x+h)=7\)

OpenStudy (anonymous):

you get \[\lim_{h\to 0}\frac{7-7}{h}=\lim_{h\to 0}0=0\]

OpenStudy (anonymous):

so what happens to the denominator?

OpenStudy (anonymous):

any fraction with zero in the numerator is zero

OpenStudy (anonymous):

you are taking a limit, not replacing \(h\) by 0 your difference quotient is identically zero for all values of \(h\neq 0\)

OpenStudy (anonymous):

so the denominator does not matter?

OpenStudy (anonymous):

you are taking a limit as the denominator goes to zero, but the whole reason for calling it "the limit" is that you cannot replace \(h\) by zero in other words when you take a limit as say \(h\to a\) you are assuming that \(h\) is close but not equal to \(a\)

OpenStudy (anonymous):

so if \(h\neq 0\) then \(\frac{0}{h}=0\) always

OpenStudy (anonymous):

and therefor the limit is zero in any case \(y=7\) is a horizontal line and as such the slope is zero for sure

OpenStudy (anonymous):

Okay, thank you, im kind of getting the jist of what your saying